# Hölder regularity of harmonic functions on metric measure spaces

### v2

@@ 4 Poisson equation on unbounded MMD spaces
Since $\mathcal{F}_{e}\ni h_{Y\cup\{x,y\}}(1_{\{x\}})-h_{\{x,y\}}(1_{\{x\}})=0$ on $\{x,y\}$ and $h_{\{x,y\}}(1_{\{y\}})\in\mathcal{F}_{e}$ is harmonic in $X\backslash\{x,y\}$, by Lemma 8.6 (4), we have
@@
$h_{Y\cup\{x\}}(1_{Y})(y)\leq\frac{R(x,y)}{R(x,X\backslash B)}.$
Thus we have
$u(y)-u(x)\leq\left(\mathrm{osc}_{Y}u\right)\frac{R(x,y)}{R(x,X\backslash B)}.$
$u-u(x)=h_{Y\cup\{x\}}(u-u(x))\geq-h_{Y\cup\{x\}}\left((\mathrm{osc}_{Y}u)1_{Y}\right)=-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y}),$
which implies
$u(y)-u(x)\geq-\left(\mathrm{osc}_{Y}u\right)h_{Y\cup\{x\}}(1_{Y})(y)\geq-\left(\mathrm{osc}_{Y}u\right)\frac{R(x,y)}{R(x,X\backslash B)}.$
@@
Lemma 8.14.
([31, THEOREM 6.3, LEMMA 9.2]) Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi$) and CS($\Psi$). Then there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, we have
$\lVert u\rVert_{L^{\infty}(B)}\leq C\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
@@
$M:=\lVert u\rVert_{L^{\infty}(\frac{3}{2}B)}\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
@@
$|u(x)-u(y)|\leq 2\lVert u\rVert_{L^{\infty}(\frac{3}{2}B)}\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\lesssim\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m.$
@@
$|u(x)-u(y)|\lesssim\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(\frac{r}{2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m\asymp\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-5.90005pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.75003pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-4.25003pt}}\!\int_{2B}|u|\mathrm{d}m,$
@@
A natural question is to consider the sharpness of the Hölder exponents $\beta_{1}-\alpha_{1}$ and $\beta_{2}-\alpha_{2}$ from HR($\Phi,\Psi$). Indeed, on a class of unbounded p.c.f. self-similar sets (see [54]) and the corresponding cable systems, these exponents are sharp, see the proof of [25, Proposition 4.1, Proposition 4.3] for the Vicsek and the SG cable systems. However, on the unbounded SC and the SC cable system, it was conjectured in [1, Conjecture 5.4, Open Question 5.5] that these exponents are not sharp, and a new fractal dimension, that is, the so-called topological Hausdorff dimension (see [4]), should be involved.
@@ 9 Some examples related to the Sierpiński carpet and the Vicsek set
In this section, we give some examples related to the Sierpiński carpet and the Vicsek set including the blowups of a fractal by another different fractal. In $\mathbb{R}^{2}$, let
@@
Then the Sierpiński carpet is the unique non-empty compact set $K$ in $\mathbb{R}^{2}$ satisfying $K=\cup_{i=1}^{8}g_{i}(K)$, see Figure 2. Let $\alpha=\log 8/\log 3$. By [42, Theorem 1.5.7, Proposition 1.5.8], the Hausdorff dimension of $K$ is $\alpha$ and the $\alpha$-dimensional Hausdorff measure $\mathcal{H}^{\alpha}(K)\in(0,+\infty)$. The construction of Brownian motion on the Sierpiński carpet was given by [6] and [46] using approximation of stochastic processes. See also [33] for another construction using $\Gamma$-convergence of quadratic forms, [9] for the construction on higher dimensional Sierpiński carpets, [17] for the construction on unconstrained Sierpiński carpets. Let $\beta=\log(8\rho)/\log 3\approx 2.09697$, where $\rho\approx 1.25147\in[\frac{7}{6},\frac{3}{2}]$ is some parameter from resistance estimates in [7, 11].
@@
Firstly, we consider the unbounded Sierpiński carpet $K^{\infty}=\cup_{n\geq 0}3^{n}K$ as follows, where $\{3^{n}K\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$. Let $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. It is obvious that V($\Phi$) holds. It was proved in [8, Theorem 1.1] that HK($\Psi$) and HHK($\Phi,\Psi$) hold. In a general setting of fractional diffusions introduced in [5, Section 3], it was proved in [1, Theorem 3.7, Lemma 3.4] that wBE($\Phi,\Psi$) and HHK($\Phi,\Psi$) hold. It was proved in [8, Theorem 8.3] that all functions in the domain of the generator of the Dirichlet form are $(\beta-\alpha)$-Hölder continuous. By the main result, Theorem 2.1, we now have the Hölder regularity for harmonic functions as HR($\Phi,\Psi$) and the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel.
Secondly, we consider the Sierpiński carpet cable system. Let $V_{0}=\{p_{1},\ldots,p_{8}\}$ and $V_{n+1}=\cup_{i=1}^{8}g_{i}(V_{n})$ for any $n\geq 0$, then $\{V_{n}\}_{n\geq 0}$ is an increasing sequence of finite subsets of $K$ and the closure of $\cup_{n\geq 0}V_{n}$ with respect to the Euclidean topology is $K$. For any $n\geq 0$, let $V^{(n)}=3^{n}V_{n}$, see Figure 4 for $V^{(0)}$, $V^{(1)}$ and $V^{(2)}$.
@@
then $(V,E)$ is an infinite, locally bounded, connected (undirected) graph, whose corresponding cable system is the Sierpiński carpet cable system, see [25, Section 3] for more detailed construction. Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$. It is obvious that V($\Phi$) holds. Using the stability result [10] for heat kernel estimates or parabolic Harnack inequalities, we have HK($\Psi$) also holds, which is similar to the case on the unbounded Sierpiński carpet. By the main result, Theorem 2.1, we have HHK($\Phi,\Psi$), $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, wBE($\Phi,\Psi$) and HR($\Phi,\Psi$) all hold. In particular, the validity of HR($\Phi,\Psi$) implies the validity of GRH($\Phi,\Psi$), which was left open in [25]. Moreover, we also have GHK($\Phi,\Psi$), which can also be given directly by $\text{HHK}_{\text{exp}}(\Phi,\Psi)$.
Thirdly, we consider the blowups of fractals related to the Sierpiński carpet and the Vicsek set. However, we blowup a fractal by another different fractal. We consider the setting in $\mathbb{R}^{2}$.
Let $K,M$ be one of the following fractals in $\mathbb{R}^{2}$: the unit interval $[0,1]\times\{0\}$, the Vicsek set and the Sierpiński carpet. Let $\alpha_{1}$ and $\beta_{1}$ be the Hausdorff dimension and walk dimension of $K$, $\alpha_{2}$ and $\beta_{2}$ the Hausdorff dimension and walk dimension of $M$. Here, $K$ serves as the “cell” fractal and $M$ provides the “model” to blowup. Let $(\mathcal{E}^{K},\mathcal{F}^{K})$ be the strongly local regular Dirichlet form on $L^{2}(K;\mu)$ with two-sided sub-Gaussian heat kernel estimates with walk dimension $\beta_{1}$, where $\mu$ is the normalized $\alpha_{1}$-Hausdorff measure on $K$, see [48, 42, 53] for the case $K$ is the Vicsek set.
@@
then $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$. Obviously, $(X,d,m,\mathcal{E},\mathcal{F})$ is an unbounded MMD space satisfying V($\Phi$). Since two-sided heat kernel estimates hold on $K,M$, by [30, Theorem 3.14], we have the elliptic Harnack inequalities and the resistance estimates both hold on $K,M$. Combining these two conditions on $K,M$, we obtain easily the corresponding elliptic Harnack inequalities and the resistance estimates on $X$. Then by [30, Theorem 3.14] again, we have HK($\Psi$) holds on $X$. Indeed, in the strongly recurrent setting, only resistance estimates are enough for the above argument, see [12, 45, 37]. By the main result, Theorem 2.1, we have HHK($\Phi,\Psi$), $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, wBE($\Phi,\Psi$) and HR($\Phi,\Psi$) all hold.
@@
    Let $S_{0}=[(0,0),(1,1)]\cup[(0,1),(1,0)]$ be the union of the diagonals of the unit square $[0,1]^{2}\subseteq\mathbb{R}^{2}$. For any $n\geq 0$, let $S_{n+1}=\cup_{i=1}^{N^{(1)}}g_{i}^{(1)}(S_{n})$. Then $\{S_{n}\}_{n\geq 0}$ is an increasing sequence of subsets of $K$ and the closure of $\cup_{n\geq 0}S_{n}$ is $K$. Let $\mathcal{S}_{0}=\cup_{n\geq 0}S_{n}$, $\nu^{(0)}$ the measure on $\mathcal{S}_{0}$ such that for any $n\geq 0$, $\nu^{(0)}|_{S_{n}}$ is the normalized 1-dimensional Hausdorff measure satisfying
@@
    For any $n\geq 0$, let $\mathcal{S}_{n+1}=\cup_{i=1}^{N^{(2)}}g_{i}^{(2)}(\mathcal{S}_{n})$, $\mathcal{S}^{(n)}=(L^{(2)})^{n}\mathcal{S}_{n}$, then similar to $\{K^{(n)}\}_{n\geq 0}$, $\{\mathcal{S}^{(n)}\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$, let $\mathcal{S}=\cup_{n\geq 0}\mathcal{S}^{(n)}$. For any cell $W\subseteq X$, we have $\mathcal{S}_{W}:=\mathcal{S}\cap W$ is a translation of $\mathcal{S}^{(0)}=\mathcal{S}_{0}$. For any two cells $W_{1},W_{2}\subseteq X$, we have either the cardinal #$\#(\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}})\in\{0,1\}$ or $\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}}$ is countably infinite, hence there exists a unique $\nu$ on $\mathcal{S}$ such that for any cell $W$, $\nu|_{\mathcal{S}_{W}}$ is the push forward of $\nu^{(0)}$ under the translation from $\mathcal{S}^{(0)}=\mathcal{S}_{0}$ to $\mathcal{S}_{W}$. Note that $\nu(B(x_{0}),r)=+\infty$ for any ball $B(x_{0},r)$ and $\nu\perp m$, that is, $\nu$ and $m$ are mutually singular.
@@
    $\leq\begin{cases}\frac{C_{3}}{t}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta_{1}}}\right)^{\frac{\beta_{1}}{\beta_{1}-1}}\right),&\text{ if }t<d(x,y),\\ \frac{C_{3}}{t}\exp\left(-C_{4}\left(\frac{d(x,y)}{t^{1/\beta_{2}}}\right)^{\frac{\beta_{2}}{\beta_{2}-1}}\right),&\text{ if }t\geq d(x,y).\\ \end{cases}$ (9.3)
    Note that $\partial p_{t}(x,\cdot)$ exists and has the above estimate $\nu$-a.e. instead of $m$-a.e.. At small scale, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scale, $\beta_{2}\approx 2.09697$, the gradient estimate has the sub-Gaussian behavior corresponding to the Sierpiński carpet. See [15, Section 3] for the results about Lipschitz and gradient estimates for heat kernel on unbounded Vicsek set, that is, the blowup of the Vicsek set by the Vicsek set itself.
@@
    Similar to the blowup of the Vicsek set by the Sierpiński carpet, we also have the Lipschitz estimate, Equation (9.2) and the gradient estimate, Equation (9.3). However, at small scale, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scale, $\beta_{2}=2$, the gradient estimate has the Gaussian behavior corresponding to the unit interval.

### v3

@@ 4 Poisson equation on unbounded MMD spaces
[…]
@@
[…]
@@
The alternative proof relies on the following Lipschitz continuity result for Green functions in the setting of Kigami’s resistance form theory.
Lemma 8.14 ([46, THEOREM 4.1]).
[…]
$\lvert g_{D}(x,y)-g_{D}(x,z)\rvert\leq R(y,z).$
An alternative proof of Proposition 8.12 using Lemma 8.14.
[…]
$\mathcal{E}(h,v)=v(x)\mathcal{E}(h,h)=\frac{v(x)}{R(x,X\backslash B)}\text{ for any }v\in\mathcal{F}_{e,B}.$
[…]
$h_{Y\cup\{x\}}(1_{Y})=1-h=\frac{g_{B}(x,x)-g_{B}(x,\cdot)}{R(x,X\backslash B)}\leq\frac{R(x,\cdot)}{R(x,X\backslash B)}.$
The remainder of the argument proceeds as in the previous proof. ∎
@@
Lemma 8.15 ([33, THEOREM 6.3, LEMMA 9.2]).
Let $(X,d,m,\mathcal{E},\mathcal{F})$ be an unbounded MMD space satisfying VD, FK($\Psi$) and CS($\Psi$). Then there exists $C\in(0,+\infty)$ such that for any ball $B=B(x_{0},r)$, for any $u\in\mathcal{F}$ which is harmonic in $2B$, we have
$\lVert{u}\rVert_{L^{\infty}(B)}\leq C\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.98003pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.26338pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.6363pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.45924pt}}\!\int_{2B}|u|\mathrm{d}m.$
@@
$M:=\lVert{u}\rVert_{L^{\infty}(\frac{3}{2}B)}\lesssim\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$
@@
$|u(x)-u(y)|\leq 2\lVert{u}\rVert_{L^{\infty}(\frac{3}{2}B)}\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\lesssim\frac{R(x,y)}{R(x,X\backslash B(x,\frac{r}{2}))}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m.$
@@
$|u(x)-u(y)|\lesssim\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(\frac{r}{2})}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m\asymp\frac{\left(\frac{\Psi}{\Phi}\right)(d(x,y))}{\left(\frac{\Psi}{\Phi}\right)(r)}\mathchoice{{\vbox{\hbox{$\textstyle-$ }}\kern-7.83337pt}}{{\vbox{\hbox{$\scriptstyle-$ }}\kern-6.11674pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.48965pt}}{{\vbox{\hbox{$\scriptscriptstyle-$ }}\kern-5.31259pt}}\!\int_{2B}|u|\mathrm{d}m,$
@@
A natural question concerns the sharpness of the Hölder exponents $\beta_{1}-\alpha_{1}$ and $\beta_{2}-\alpha_{2}$ arising from HR($\Phi,\Psi$). Indeed, on a class of unbounded p.c.f. self-similar sets and the corresponding cable systems, these exponents are sharp, see the proof of [26, Proposition 4.1, Proposition 4.3] for the Vicsek and the Sierpiński gasket cable systems. However, on the unbounded Sierpiński carpet and the Sierpiński carpet cable system, it was conjectured in [2, Conjecture 5.4, Open Question 5.5] that these exponents are not sharp, and a new fractal dimension, the so-called topological Hausdorff dimension (see [5]), should be involved.
@@ 9 Some examples related to the Sierpiński carpet and the Vicsek set
In this section, we give some examples related to the Sierpiński carpet and the Vicsek set, including blowups of one fractal by another. In $\mathbb{R}^{2}$, let
@@
Then the Sierpiński carpet is the unique non-empty compact set $K$ in $\mathbb{R}^{2}$ satisfying $K=\cup_{i=1}^{8}g_{i}(K)$, see Figure 2. Let $\alpha=\log 8/\log 3$. By [44, Theorem 1.5.7, Proposition 1.5.8], the Hausdorff dimension of $K$ is $\alpha$ and the $\alpha$-dimensional Hausdorff measure $\mathcal{H}^{\alpha}(K)\in(0,+\infty)$. The construction of Brownian motion on the Sierpiński carpet was given by [7] and [48] using approximations of stochastic processes. See also [35] for another construction using $\Gamma$-convergence of quadratic forms, [10] for the construction on higher dimensional Sierpiński carpets, [18] for the construction on unconstrained Sierpiński carpets. Let $\beta=\log(8\rho)/\log 3\approx 2.09697$, where $\rho\approx 1.25147\in[\frac{7}{6},\frac{3}{2}]$ is some parameter from resistance estimates in [8, 12].
@@
Firstly, we consider the unbounded Sierpiński carpet $K^{\infty}=\cup_{n\geq 0}3^{n}K$ as follows, where $\{3^{n}K\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$. Let $\alpha_{1}=\alpha_{2}=\alpha$, $\beta_{1}=\beta_{2}=\beta$. It is obvious that V($\Phi$) holds. It was proved in [9, Theorem 1.1] that HK($\Psi$) and HHK($\Phi,\Psi$) hold. In a general setting of fractional diffusions introduced in [6, Section 3], it was proved in [2, Theorem 3.7, Lemma 3.4] that wBE($\Phi,\Psi$) and HHK($\Phi,\Psi$) hold. It was proved in [9, Theorem 8.3] that all functions in the domain of the generator of the Dirichlet form are $(\beta-\alpha)$-Hölder continuous. By the main result, Theorem 2.1, we now have the Hölder regularity for harmonic functions HR($\Phi,\Psi$) and the Hölder estimate $\text{HHK}_{\text{exp}}(\Phi,\Psi)$ with exponential terms for the heat kernel.
Secondly, we consider the Sierpiński carpet cable system. Let $V_{0}=\{p_{1},\ldots,p_{8}\}$ and $V_{n+1}=\cup_{i=1}^{8}g_{i}(V_{n})$ for any $n\geq 0$, then $\{V_{n}\}_{n\geq 0}$ is an increasing sequence of finite subsets of $K$, and the closure of $\cup_{n\geq 0}V_{n}$ with respect to the Euclidean topology is $K$. For any $n\geq 0$, let $V^{(n)}=3^{n}V_{n}$, see Figure 4 for $V^{(0)}$, $V^{(1)}$ and $V^{(2)}$.
@@
then $(V,E)$ is an infinite, locally bounded, connected (undirected) graph, whose corresponding cable system is the Sierpiński carpet cable system, see [26, Section 3] for a more detailed construction. Let $\alpha_{1}=1$, $\beta_{1}=2$, $\alpha_{2}=\alpha$, $\beta_{2}=\beta$. It is obvious that V($\Phi$) holds. Using the stability result [11] for heat kernel estimates or parabolic Harnack inequalities, we have HK($\Psi$) also holds, which is similar to the case on the unbounded Sierpiński carpet. By the main result, Theorem 2.1, we have HHK($\Phi,\Psi$), $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, wBE($\Phi,\Psi$) and HR($\Phi,\Psi$) all hold. In particular, the validity of HR($\Phi,\Psi$) implies the validity of GRH($\Phi,\Psi$), which was left open in [26]. Moreover, we also have GHK($\Phi,\Psi$), which can also be given directly by $\text{HHK}_{\text{exp}}(\Phi,\Psi)$.
Thirdly, we consider blowups of fractals related to the Sierpiński carpet and the Vicsek set. However, we blowup one fractal by a different fractal. In what follows, we work in $\mathbb{R}^{2}$.
Let $K,M$ be one of the following fractals in $\mathbb{R}^{2}$: the unit interval $0,1]\times\{0\}$, the Vicsek set, and the Sierpiński carpet. Let $\alpha_{1}$ and $\beta_{1}$ be the Hausdorff dimension and the walk dimension of $K$, and let $\alpha_{2}$ and $\beta_{2}$ be the Hausdorff dimension and the walk dimension of $M$. Here, $K$ serves as the “cell” fractal and $M$ provides the “model” to blowup. Let $(\mathcal{E}^{K},\mathcal{F}^{K})$ be the strongly local regular Dirichlet form on $L^{2}(K;\mu)$ with two-sided sub-Gaussian heat kernel estimates with walk dimension $\beta_{1}$, where $\mu$ is the normalized $\alpha_{1}$-Hausdorff measure on $K$, see [50, 44, 57] for the case $K$ is the Vicsek set.
@@
then $(\mathcal{E},\mathcal{F})$ is a strongly local regular Dirichlet form on $L^{2}(X;m)$. Obviously, $(X,d,m,\mathcal{E},\mathcal{F})$ is an unbounded MMD space satisfying V($\Phi$). Since two-sided heat kernel estimates hold on $K,M$, by [31, Theorem 3.14], we have the elliptic Harnack inequalities and the resistance estimates both hold on $K,M$. Combining these two conditions on $K,M$, we obtain easily the corresponding elliptic Harnack inequalities and the resistance estimates on $X$. Then by [31, Theorem 3.14] again, we have HK($\Psi$) holds on $X$. Indeed, in the strongly recurrent setting, only resistance estimates are enough for the above argument, see [13, 47, 39]. By the main result, Theorem 2.1, we have HHK($\Phi,\Psi$), $\text{HHK}_{\text{exp}}(\Phi,\Psi)$, wBE($\Phi,\Psi$) and HR($\Phi,\Psi$) all hold.
@@
    Let $S_{0}=[(0,0),(1,1)]\cup[(0,1),(1,0)]$ be the union of the diagonals of the unit square $0,1]^{2}\subseteq\mathbb{R}^{2}$. For any $n\geq 0$, let $S_{n+1}=\cup_{i=1}^{N^{(1)}}g_{i}^{(1)}(S_{n})$. Then $\{S_{n}\}_{n\geq 0}$ is an increasing sequence of subsets of $K$, and the closure of $\cup_{n\geq 0}S_{n}$ is $K$. Let $\mathcal{S}_{0}=\cup_{n\geq 0}S_{n}$, $\nu^{(0)}$ the measure on $\mathcal{S}_{0}$ such that for any $n\geq 0$, $\nu^{(0)}|_{S_{n}}$ is the normalized 1-dimensional Hausdorff measure satisfying
@@
    For any $n\geq 0$, let $\mathcal{S}_{n+1}=\cup_{i=1}^{N^{(2)}}g_{i}^{(2)}(\mathcal{S}_{n})$, $\mathcal{S}^{(n)}=(L^{(2)})^{n}\mathcal{S}_{n}$, then similar to $\{K^{(n)}\}_{n\geq 0}$, $\{\mathcal{S}^{(n)}\}_{n\geq 0}$ is an increasing sequence of subsets of $\mathbb{R}^{2}$, let $\mathcal{S}=\cup_{n\geq 0}\mathcal{S}^{(n)}$. For any cell $W\subseteq X$, we have $\mathcal{S}_{W}:=\mathcal{S}\cap W$ is a translation of $\mathcal{S}^{(0)}=\mathcal{S}_{0}$. For any two cells $W_{1},W_{2}\subseteq X$, we have either the cardinal $(\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}})\in\{0,1\}$ or $\mathcal{S}_{W_{1}}\cap\mathcal{S}_{W_{2}}$ is countably infinite, hence there exists a unique $\nu$ on $\mathcal{S}$ such that for any cell $W$, $\nu|_{\mathcal{S}_{W}}$ is the push forward of $\nu^{(0)}$ under the translation from $\mathcal{S}^{(0)}=\mathcal{S}_{0}$ to $\mathcal{S}_{W}$. Note that $\nu(B(x_{0}),r)=+\infty$ for any ball $B(x_{0},r)$ and $\nu\perp m$, that is, $\nu$ and $m$ are mutually singular.
@@
[…]
    Note that $\partial p_{t}(x,\cdot)$ exists and has the above estimate $\nu$-a.e. instead of $m$-a.e.. At small scales, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scales, $\beta_{2}\approx 2.09697$, the gradient estimate has the sub-Gaussian behavior corresponding to the Sierpiński carpet. See [16, Section 3] for the results about Lipschitz and gradient estimates for heat kernel on unbounded Vicsek set, that is, the blowup of the Vicsek set by the Vicsek set itself.
@@
    Similar to the blowup of the Vicsek set by the Sierpiński carpet, we also have the Lipschitz estimate, Equation (9.2) and the gradient estimate, Equation (9.3). However, at small scales, $\beta_{1}=\log 15/\log 3$, the gradient estimate has the sub-Gaussian behavior corresponding to the Vicsek set, while at large scales, $\beta_{2}=2$, the gradient estimate has the Gaussian behavior corresponding to the unit interval.